OpenStudy (anonymous):

the following is true or false where m, n are natural numbers n^2=m^2+1002 Explain how?

OpenStudy (unklerhaukus):

\[n^2=m^2+1002\]\[n^2-m^2=1002\]\[(n+m)(n-m)=1002\]

OpenStudy (unklerhaukus):

\[1002=2\times501=2\times3\times167\]

OpenStudy (anonymous):

what are the examples of combination of m and n?

OpenStudy (unklerhaukus):

i can not think of any that work